### v1

@@ 2.1 Notation
We denote a neural network as $\mathbb{F}$, its approximation as $\widetilde{\mathbb{F}}$, benign inputs and label as ($x$, $y$). Perturbations for adversarial samples are represented by $\delta$. A small value is represented by $\epsilon$. The step length for PGD attacks is represented by $\alpha$. The precision for PANN is $2^{-\beta}$ determined by $\beta$. The approximation of $\operatorname{sgn}$ and $\operatorname{ReLU}$ are denoted by $\operatorname{Appsgn}$ and $\operatorname{AppReLU}$ respectively.
@@ 2.2 Polynomial Approximation of Neural Networks
Common approach for approximation in PANNs include simple polynomials, Taylor polynomials, Minimax approximation, etc. Among them, Minimax approximation is the most common because it can approximate for an interval and is computationally efficient. Therefore, our focus is primarily on studying PANN with Minimax approximation.
@@
This paper refers to $p(z)$ as a polynomial approximation function. An $\beta$-close polynomial approximation of a function $f(z)$ in interval $[a,b]$ satisfies:
$|p(z)-f(z)|\leq 2^{-\beta},\quad x\in[a,b]$ (1)
For many neural networks, activation functions, especially the ReLU function, are the main part requiring approximation and naturally is the main cause of approximation errors. ReLu is typically approximated using the formula $\operatorname{ReLU}(z)=\frac{z+z\operatorname{sgn}(z)}{2}$. If the error for $\operatorname{Appsgn}(z)$ is $\epsilon_{sgn}$, then the error for ReLU approximation $\operatorname{AppReLU}(z)=\frac{z+z\operatorname{Appsgn}(z)}{2}$ is:
$\epsilon_{relu}:=\frac{z+z\operatorname{Appsgn}(z)}{2}-\operatorname{ReLU}(z)=\frac{\epsilon_{sgn}\cdot z}{2}$ (2)
If $\operatorname{Appsgn}(z)$ is $\beta$-close, the approximation of $\operatorname{ReLu}(z)$ will have error bound $2^{-(\beta+1)}z$. Equation 2 shows an important fact that the larger the input $z$, the larger the approximation error $\epsilon_{relu}$.
It is also worth noting that $p(z)$ is composed by small degree polynomials: $p=p_{k}\circ p_{k-1}\circ\dots\circ p_{1}$. These small degree polynomials are chosen according to precision. Therefore, PANN with different precision uses different polynomial approximations, leading to different error locations. Therefore, it is hard to improve PANN across different precision simultaneously via fune-tuning.
3 Effects of Approximation Errors on PANN
This section presents our findings regarding approximation errors on neural networks, highlighting its similarities and differences with perturbations in adversarial samples (detailed in Appendix B), and also how neural networks are affected by weight regularization when approximation errors ¡0are involved.
3.1 Perturbations on Different Information
@@
Figure 2: Minimax approximation errors for $\operatorname{Appsgn}(z)$ on $[-1,-\epsilon]$ with precision $2^{-10}$ (left) and $2^{-14}$ (right)
Activation functions are the main components requiring approximation in existing PANNs. Thus, we focus our analyses on the error introduced by ReLU, the most common activation function. As discussed in Section 2.2, approximation of ReLU, $\operatorname{AppReLu}$, is computed by $\frac{z+z\operatorname{Appsgn}(z)}{2}$.The approximation error of $\operatorname{Appsgn}$ function in $[-1,-\epsilon]$ is shown in Figure 2. We note that since $\operatorname{Appsgn}(z)$ is odd, the same approximation errors will also be observed on $[1,\epsilon]$. Assume the error for approximating $\operatorname{sgn}(z)$ is $\epsilon_{sgn}\in[-2^{-\beta},2^{\beta}]$ , we have:
$\epsilon_{relu}:=\operatorname{AppReLU}(z)-\operatorname{ReLU}(z)=\frac{\epsilon_{sgn}\cdot z}{2}$ (3)
We discuss a subtle issue regarding approximation error in PANN involving ReLU. In the backbone model $\operatorname{ReLU(z)}=0$ when $z<0$. In other words, perturbations added to input with $z<0$ will not affect the next layer, whereas approximation error introduced with $z<0$ will be propagated to the next layer. In more detail, we denote the inputs to $\operatorname{ReLU}$ and $\operatorname{AppReLU}$ in a neural network as $z=\boldsymbol{W}_{i}x$. We consider the following two cases: $\boldsymbol{W}_{i}x>0$ and $\boldsymbol{W}_{i}x<0$.
When $\boldsymbol{W}_{i}x>0$, $\operatorname{ReLU}(\boldsymbol{W}_{i}x)=\boldsymbol{W}_{i}x$ in the backbone model, the information of $x$ will be passed to the next layer and may contribute to the final results. We call this “information contributing to the outputs”. For this information, $\epsilon_{relu}=\frac{\epsilon_{sgn}\cdot\boldsymbol{W}_{i}x}{2}$. Since neural networks are nearly linear within a very small interval, we have:
$\operatorname{AppReLU}(\boldsymbol{W}_{i}x)$ $=(1+\frac{\epsilon_{sgn}}{2})\boldsymbol{W}_{i}x$ (4)
$=\operatorname{ReLU}(\boldsymbol{W}_{i}(x+\frac{\epsilon_{sgn}}{2}x))$
If these approximation errors lead to a mistake in the output, their effect is the same as an adversarial sample with perturbation $\frac{\epsilon_{sgn}}{2}x$ on the inputs $x$.
When $\boldsymbol{W}_{i}x<0$, $\operatorname{ReLU}(\boldsymbol{W}_{i}x)=0$ always holds. The information will not be passed to the next layer and contribute to outputs, which we mention as “irrelevant information in the input background”. For this information, we have:
$\operatorname{AppReLU}(\boldsymbol{W}_{i}x)=\frac{\epsilon_{sgn}}{2}\boldsymbol{W}_{i}x=\epsilon_{relu}$ (5)
Different from the previous case, if these errors lead to a wrong output, they can hardly be related to adversarial samples. Since perturbations in adversarial samples are very small and $z$ may be far from 00, it is difficult to make such a large change in $z$ by a finite input perturbation.
In short, the effects of approximation errors in the former case is similar to an evasion attack. In contrast, the second case is unique to the approximation errors in PANNs. We note one more difference, approximation errors differ from adversarial samples as they do not follow the directions of gradients, meaning that while some errors are harmful, some can be beneficial. Nonetheless, approximation error in the second case is most likely harmful to the PANNs because as long as $\boldsymbol{W}_{i}x\neq 0$, errors will lead to unknown behavior in neural networks.
3.2 Observation through Input Perturbations
As discussed, approximation errors affect information contributing to the outputs and irrelevant information in the input background. The effects of approximation errors in the former case are similar to an evasion attack and provide an explanation on how they affect inference accuracy in PANNs. However, it is unclear if approximation errors in the latter case also affect PANNs’ accuracy.
To further investigate the issue, we make use of the fact that $\epsilon_{relu}$ contains the product $\frac{\epsilon_{sgn}\cdot\boldsymbol{W}_{i}x}{2}$. More concretely, we can adjust the magnitude of $\epsilon_{relu}$ by controlling the magnitude of $\boldsymbol{W}_{i}x$. We can seperate the effect of the two cases by adding perturbations to $x$ with $\boldsymbol{W}_{i}x>0$ or $\boldsymbol{W}_{i}x<0$.
For information contributing to the outputs, we can add perturbations to the part of $x$ with $\boldsymbol{W}_{i}x>0$, which is usually the main part of inputs (e.g., the objects in images). If there is an adversarial sample $x+\delta_{app}$ for PANN, we get:
$\operatorname{AppReLU}(\boldsymbol{W}_{i}(x+\delta_{app}))$ (6)
$=(1+\frac{\epsilon_{sgn}}{2})\boldsymbol{W}_{i}(x+\delta_{app})$
$=\operatorname{ReLU}(\boldsymbol{W}_{i}(x+\delta_{app}+\frac{\epsilon_{sgn}}{2}(x+\delta_{app}))$
It means if a valid input perturbation $\delta_{app}$ for PANN exists, we may find a $\delta_{bb}=\delta_{app}+\frac{\epsilon_{sgn}}{2}(x+\delta_{app})$ for the backbone model. Since $\delta_{app}$ has driven $x$ to its class margin and $\frac{\epsilon_{sgn}}{2}$ is very small, $\delta_{bb}$ is easy to find by slightly enlarging $\delta_{app}$.
For irrelevant information in the input background, we can add perturbations to the part of $x$ with $\boldsymbol{W}_{i}x<0$. This part is generally the background of inputs. Different from the previous case, if we find a valid $x_{adv}=x+\delta_{app}$ for PANN here, we have:
$\operatorname{AppReLU}(\boldsymbol{W}_{i}(x+\delta_{app}))=\boldsymbol{W}_{i}\frac{\epsilon_{sgn}}{2}(x+\delta_{app})$ (7)
This error is unreachable by $\operatorname{ReLU}(\boldsymbol{W}_{i}(x+\delta_{bb}))$ for a small $\delta_{bb}$, because $\operatorname{ReLU}$ always outputs 0 for negative inputs. If $\operatorname{AppReLU}(\boldsymbol{W}_{i}(x+\delta_{app}))>0$, we need $\operatorname{ReLU}(\boldsymbol{W}_{i}(x+\delta_{bb}))>0$ then $\boldsymbol{W}_{i}\delta_{bb}>-\boldsymbol{W}_{i}x$, which may require a large $|\delta_{bb}|$; Moreover, if $\operatorname{AppReLU}(\boldsymbol{W}_{i}(x+\delta_{app}))<0$, it’s impossible to find a $\delta_{bb}$ such that $\operatorname{ReLU}(\boldsymbol{W}_{i}(x+\delta_{bb}))<0$. If $\delta_{app}$ can easily mislead PANN, we can identify that approximation errors on irrelevant information in the input background can affect PANN significantly. Meanwhile, if we can hardly find $\delta_{app}$ works on the backbone, we can conclude that this information is irrelevant to the task neural networks are trained for, and that crafting perturbations to this part during training has little damage to the accuracy of the backbone model.
3.3 Effects of weight regularization on PANN
Figure 3: Backbone models (bb) trained with different weight decay (wd) have similar accuracy (dot line). However, large weight decay can reduce their accuracy on PANN as the epoch increases. (ResNet-20, CIFAR-10, precision $2^{-8}$)
Weight regularization is crucial in model training. It can improve model generalization and accuracy Zhang et al. (2018a). However, we observed that it is very harmful to PANN. Even a small weight regularization can significantly reduce the “sturdiness” of the neural networks. An example of PANN on models trained with different weight decay is presented in Figure 3. Their backbone (bb) models have similar accuracy, but the performances of their corresponding PANNs vary greatly. In the early training stages, all these models have a solid resistance to approximation errors, but this resistance is destroyed by weight regularization as the epochs increase. Larger weight decay can make this reduction quicker. At epoch 200, the model with weight decay 0 only reduced the accuracy by about 10%, but weight decay 0.001 reduced the accuracy by over 80%. This reduction is observed in many model structures and has become an important factor affecting the accuracy of PANN (shown in Table 2 and Table 3).
We attribute this to the fact that weight regularization can make the landscape of neural nets more complicated Ziyin et al. (2022). This complexity may not be problematic for normal models, but it can be detrimental to PANN when combined with approximation errors. Though weight regularization can improve generalization within input distribution, approximation errors, especially the errors on irrelevant information in the input background, can be far away from this scope, rendering them useless for PANN.
@@ 4 Improving low-precision PANN’s accuracy
4.1 AT-like Solutions for Improving PANN
We propose a solution to enhance “sturdiness” based on the findings about the relationship between approximation error and adversarial robustness discussed in Section 3.1. This inspire us to adapt adversarial training to deal with approximation errors.
@@ 4.1 AT-like Solutions for Improving PANN
Training Methods: For a neural network with layered architectures, the input layer is numbered 00, the output layer is numbered $L$, and the hidden layers $h_{l}$ are numbered $1$, $\dots,L-1$. We denote $z_{l}$ as the input at layer $l$ before the activation function and $z_{L}$ is the output of the neural network, $x_{l}$ as the vector after application of the activation function (to $z_{l-1}$) and $x_{0}=x$ is the neural network input. For normal models, $z_{l}=h_{l}(x_{l})$. For our method, we use $\mathbb{F}^{\prime}(x)$ to represent the $\mathbb{F}(x)$ replacing all $z_{l}$ with $z^{\prime}_{l}=h_{l}(x_{l}+\epsilon_{in,l})$, use $\beta^{\prime}$ to represent the amplification due to the multiplication of errors with intermediate values. We are looking for a neural network that satisfies the following:
@@
$\inf_{\theta}\mathbb{E}_{\mathbb{P}}[\mathcal{L}(\mathbb{F}^{\prime}(x),y)],\quad\forall||\epsilon_{in,l}||<2^{-\beta+\beta^{\prime}}$ (8)
@@
When trying to design this AT-like method, there are two considerations. First, adversarial samples only affect information contributing to the outputs, but we analyze and prove that approximation errors can also affect irrelevant information in the input backgrounds. Second, adversarial samples only add perturbations to inputs, while approximation errors are generated in almost every layer. These differences result in a much wider range of perturbations in PANN, with a much larger amplitude (after accumulated and amplified). Therefore, we should not introduce perturbations during training at every point susceptible to approximation errors like AT does because the accuracy of the backbone model will be reduced. This reduction will limit the upper bound of PANN accuracy and affect the use of the backbone model in other areas, thereby increasing the overall costs.
To address this issue, we consider restricting perturbations only to specific positions. According to Equation 3, approximation errors are multiplied by function inputs, which means that errors on activation function inputs with large absolute values are more detrimental. Therefore, we can add perturbations only to the proportion of activation function inputs with the largest magnitude. Furthermore, intra-model perturbations consist of two parts with different characters. One part affects information contributing to the outputs and can be reached by small perturbations on the backbone model input (usually the positive inputs to ReLU), denoted by $e_{l}(x+\delta_{bb})$. The other part affects irrelevant information in the input background and is unreachable by $\delta_{bb}$ (usually the negative inputs to ReLU), donated by $\epsilon_{irr,l}$. So we have $\epsilon_{int,l}=e_{l}(x+\delta_{bb})+\epsilon_{irr,l}$. The first part does not follow the gradient direction and is not effective enough. In contrast, the second part is always harmful (explained in Section 3.1). This means that adding perturbations to negative values with large magnitudes is the most effective and has minimal effect on the backbone model’s performance.
@@
With the reasons above, we design a noise generator for negative values with large magnitude (NGNV) in activation function inputs during the training phase. For the smallest $r$ (eg: 0.7) percent negative values in the ReLU inputs, we generate Gaussian noise $\epsilon_{gau}$ to simulate $\epsilon_{sgn}$ and multiply with these values. The products are then multiplied by a parameter $\lambda$ (eg: 0.05) and added to ReLU inputs. In other words, for a ReLU input $z$, we generate intra-model perturbations $\delta_{int}$ as:

### v2

@@ 4.1 Notation
We denote a neural network as $\mathbb{F}$, benign inputs and label as ($x$, $y$). we use $\widetilde{\cdot}$ (i.e., $\widetilde{y}$) to represent the values in PANN (with approximation errors). We use $W_{i}$ to denote the $i$-th row vector of a matrix $W$. A small value is represented by $\epsilon$. The error bound for approximation is $2^{-\beta}$ determined by $\beta$. The approximation of $\operatorname{sgn}$ and $\operatorname{ReLU}$ are denoted by $\operatorname{Appsgn}$ and $\operatorname{AppReLU}$ respectively.
@@ 4.2 Polynomial Approximation of Neural Networks
Polynomial Approximation of Neural Networks (PANN) is the neural network replacing all (or part of) non-polynomial functions with polynomial approximations. Common approaches for approximation in PANNs include low-degree polynomials, Taylor polynomials, Minimax approximation, etc. Among these methods, low-degree polynomial is the fastest but yields low accuracy, thus generally requires customized model structures and fine-tuning models directly on PANN. In contrast, Minimax approximation with high-degree polynomials is slow but can yield negligible approximation errors, which can achieve PANN with the same accuracy as backbones.
@@
In this paper, we refers to $p(z)$ as a polynomial approximation of a function $g(z)$. The approximation error of this approximation is defined as:
$\epsilon_{g}(z)=p(z)-g(z)$ (13)
4.2.1 Simple approximation with low-degree polynomials
Since training directly on PANN with simple polynomials [4, 14] is not an approximation of the existing model structure and parameters (discussed in Section 3.1) and PANNs replacing all nonlinear functions are usually targeting on high-degree polynomials or precise approximation [11, 3]. We only introduce ReLU replacement schemes which replace a partial of non-polynomial functions [17, 2].
ReLU replacement aims at replacing a partial of ReLU activation functions with polynomial functions such as $p(z)=z$ [17] and $p(z)=0.14z^{2}+0.5z+0.28$ [2]. Given a pre-trained model, this process can be achieved by the following three steps: (1) Replace all activation functions (ReLU) $g(z)$ in it with:
$\sigma(z)=c\cdot g(z)+(1-c)\cdot p(z)$ (14)
(2) Train the model to reduce the value of $|c|$ by designing a proper loss function, while fine-tuning the pre-trained weights to suit the new activation functions; (3) Binarize $c$ depending on if the value $c$ satisfies some conditions (e.g.: if $c$ is smaller than a threshold value), and get the final activation function equal to $g(z)$ or $p(z)$; (4) Freeze $c$ and fine-tune the model based on the obtained activation functions.
ReLU replacement schemes will repeat the above steps until the ReLU counts are reduced to a certain number. This process results in two problems. First, the pre-trained model parameters were designed for the original structure using ReLU functions, and it was never intended to use polynomials from the beginning. Even with fine-tuning, it is difficult to achieve the same accuracy as the baseline (especially when the number of ReLUs is very small).
Second, the model weights are fine-tuned for the function $g^{\prime}(z)=c\cdot g(z)+(1-c)\cdot p(z)$, while the activation function in the final result is $g(z)$ or $p(z)$. This introduces an approximation error:
$\epsilon_{g}(z)=\left\{\begin{aligned} &(1-c)\cdot g(z)-(1-c)\cdot p(z),&\sigma(z)=g(z)\\ &-c\cdot g(z)+c\cdot p(z),&\sigma(z)=p(z)\\ \end{aligned}\right.$ (15)
These two factors together result in the accuracy reduction in the obtained PANN, and they both indicate the same issue: the model is not sufficiently resistant to approximation errors, which we call poor “sturdiness”.
4.2.2 Precise approximation with Minimax approximation
In Minimax approximation, $p(z)$ is composed by small degree polynomials: $p=p_{k}\circ p_{k-1}\circ\dots\circ p_{1}$. The degree and parameters of small-degree polynomials are chosen according to approximation precision. The approximation precision is usually determined by a parameter $\beta$, which means that the absolute value of approximation error $\epsilon_{g}(z)$ is bounded by $2^{-\beta}$: $|\epsilon_{g}(z)|<2^{-\beta}$. An $\beta$-close polynomial approximation of a function $g(z)$ satisfies:
$|p(z)-g(z)|\leq 2^{-\beta}$ (16)
@@ 4.2.2 Precise approximation with Minimax approximation
Figure 3: Approximation errors for the Minimax approximation of $\operatorname{sgn}(z)$ on $[-1,-\epsilon]$ with precision $2^{-10}$ (left) and $2^{-14}$ (right)
When used in cryptographic schemes, parameter $\beta$ (or the corresponding ) is an important indicator of computation cost. Because a large $\beta$ (high precision) requires a $p(z)$ with a large degree, which involves a large amount of multiplication in cryptography computation. This phenomenon is extremely significant in HE-based schemes since successive multiplications require bootstrapping (decryption and re-encryption in ciphertext), which is the most expensive part in HE. Therefore, reducing the effects of approximation errors and thus allowing for low precision (or degree) is an important topic.
For most PANN schemes, activation functions, especially the ReLU function, are the main part requiring approximation and naturally are the main cause of approximation errors. Since the ReLU function itself is hard to approximate in Minimax directly, it’s typically approximated by the sgn function using the formula:
$\operatorname{AppReLU}(z)=\frac{z+z\operatorname{Appsgn}(z)}{2}$ (17)
Define the error for approximating the function $\operatorname{sgn}(z)$ as:
$\epsilon_{sgn}(z)=\operatorname{Appsgn}(z)-\operatorname{sgn}(z)$ (18)
Then the approximation error for ReLU approximation is:
$\epsilon_{relu}(z)=\frac{\epsilon_{sgn}(z)z}{2}$ (19)
If the $\operatorname{Appsgn}$ satisfies $\beta$-close that $|\epsilon_{sgn}|<2^{-\beta}$, we can get:
$|\epsilon_{relu}(z)|<2^{-\beta}|z|$ (20)
Equation 20 illustrates that the magnitude of approximation errors is proportional to the magnitude of $z$ (the input to ReLU).
4.3 Mixup
Mixup is the method that introduces interpolated samples to the training set [30], which was proved to help model robustness and generalization [31]. For parameter $\lambda$ and samples $(x,y)$ and $(x^{\prime},y^{\prime})$ randomly selected from the training set, the interpolated sample is:
$\widetilde{x}=\lambda x+(1-\lambda)x^{\prime},\quad\widetilde{y}=\lambda y+(1-\lambda)y^{\prime}$ (21)
TABLE II: Accuracy of PANN on ResNet-20, Shufflenetv2, DLA-34, Mobilenetv2 trained with weight decay (wd) 0, 1e-4 and 5e-4. We give Top-1 accuracy of the backbone (baseline) and PANN with precision $2^{-\beta}$ ($2^{-8}$ and $2^{-9}$). Training methods include Vanilla, Mixup, and robustness-based solution NGNV. The results with the highest accuracy for each model and approximation precision are bolded.
Vanilla Mixup NGNV Mixup+NGNV
$\beta$ wd 0 1e-4 5e-4 0 1e-4 5e-4 0 1e-4 5e-4 0 1e-4 5e-4
[…]
Baseline 90.39 91.57 92.14 91.30 91.60 92.17 91.10 91.70 92.49 91.49 91.59 92.16
[…]
Baseline 88.54 90.22 88.60 89.08 90.19 87.62 88.66 90.69 89.60 89.10 90.54 88.32
[…]
Baseline 93.43 94.38 95.10 94.96 95.12 95.41 93.85 94.92 95.21 95.09 95.62 95.58
[…]
Baseline 92.69 93.49 91.45 93.60 93.74 90.50 92.97 93.58 91.84 93.68 94.03 91.58
@@ 5 Improving low-precision PANN’s accuracy
5.1 Negative Inputs Leakage Method
We propose a solution to enhance “sturdiness” based on Theorem 2.1 that approximation errors on the negative inputs of ReLU increase more loss than those on positive inputs. Intuitively, because some gradients stop updating due to negative inputs to ReLU having zero derivative, we can leak the information of ReLU’s negative inputs to the next layer and can guide the training to approach the stationary point of PANN, thus increasing the sturdiness.
@@ 5.1 Negative Inputs Leakage Method
Formally, for a neural network with layered architectures, the input layer is numbered 00, the output layer is numbered $H$, and the hidden layers $h_{l}$ are numbered $1$, $\dots,H-1$. We denote $z_{l}$ as the input at layer $l$ before the activation function and $z_{H}$ is the output of the neural network, $x_{l}$ as the vector after application of the activation function (to $z_{l-1}$) and $x_{0}=x$ is the neural network input. For normal models, $z_{l}=h_{l}(x_{l})$ and $x_{l+1}=\operatorname{ReLU}(z_{l})$. For our method, we use $\overline{\mathbb{F}}(x)$ to represent the $\mathbb{F}(x)$ replacing all $x_{l+1}$ with $\overline{x}_{l+1}=\operatorname{z_{l}}+\epsilon_{in,l}$, use $\beta^{\prime}$ to represent the amplification due to the multiplication of errors with intermediate values. We are looking for a neural network that satisfies the following:
@@
$\inf_{\theta}\mathbb{E}_{\mathbb{P}}[L(\overline{\mathbb{F}}(x),y)],\quad\forall||\epsilon_{in,l}||<2^{-\beta+\beta^{\prime}}$ (22)
@@
When trying to implement this method, there are two considerations. First, approximation errors affect every input to ReLU. Second, approximation errors are generated in almost every layer. The wide-range perturbations after accumulated and amplified result in a much larger noise amplitude, which will affect the training and reduce the accuracy of the backbone model (baseline), which can limit the upper bound of PANN accuracy.
Therefore, we consider restricting perturbations only to specific positions by taking into account two specific properties. First, as implied in Theorem 2.1, approximation errors on ReLU’s negative inputs are more harmful. Second, approximation errors are usually the polynomial of ReLU inputs $z$ such that the error amplitude is proportional to the amplitude of $z$. The two properties suggest that adding perturbations to negative values with large magnitudes is the most effective and has minimal effect on the backbone model’s performance.
@@
With the reasons above, we design a noise generator for negative values with large magnitude (NGNV) in activation function inputs during the training phase. For the smallest $r$ (e.g.: 0.3) percent negative values in the ReLU inputs (the negative value with the largest magnitudes), we generate Gaussian noise $\epsilon_{gau}$ to simulate $\epsilon_{sgn}$ and multiply with these values. The products are then multiplied by a parameter $\lambda$ (e.g.: 0.05) and added to ReLU inputs. In other words, for a ReLU input $z$, we generate intra-model perturbations $\delta_{int}$ as: